OpenStudy (anonymous):

find the solution to -x^2-2x-5=0. medal for best answer

OpenStudy (anonymous):

@abb0t u know how to do this

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@hero alittle help

hero (hero):

-x^2-2x-5=0 -(x^2 + 2x + 5) = 0 x^2 + 2x + 5 = 0 b^2 - 4ac = 2^2 - 4(1)(5) = 4 - 20 = -16 = no real solutions

OpenStudy (anonymous):

appreciate it hero